The total book value of WTC’s equity is $10 million, and book value is $20 per share. The stock market has a market-to-book ratio of 1.5 and cost of equity is 15%. The firm’s bonds have a face value of $5 million and sell at a price 110% of face value. The yield to maturity on the bonds is 9%, and the firm’s tax rate is 21%. Find the company’s WACC.
Answer #1

WACC = Weight of debt * Cost of Debt * (1 - Tax) + Weight of equity * Cost of Equity

Weight of debt and equity is based on market value of firm.

Market Value of Equity = Book Value of Equity * Market to book ratio = $10 million * 1.5 = $15 million.

Market value of debt = Face Value * % of par value = $5 million * 110% = $5.50 million

Weight of debt = Debt/(Debt + Equity) = $5.50/($5.50 + $15) = 26.83%

Weight of equity = Equity/(Debt + Equity) = $150/($5.50 + $15) = 73.17%

WACC = 26.83% * 9% * (1 - 21%) + 73.17% * 15%

WACC = 12.88%